# Citation-Enhanced Generation for LLM-based Chatbots

### v3

@@ Abstract
Large language models (LLMs) exhibit powerful general intelligence across diverse scenarios, including their integration into chatbots. However, a vital challenge of LLM-based chatbots is that they may produce hallucinated content in responses, which significantly limits their applicability. Various efforts have been made to alleviate hallucination, such as retrieval augmented generation and reinforcement learning with human feedback, but most of them require additional training and data annotation. In this paper, we propose a novel post-hoc Citation-Enhanced Generation (CEG) approach combined with retrieval argumentation. Unlike previous studies that focus on preventing hallucinations during generation, our method addresses this issue in a post-hoc way. It incorporates a retrieval module to search for supporting documents relevant to the generated content, and employs a natural language inference-based citation generation module. Once the statements in the generated content lack of reference, our model can regenerate responses until all statements are supported by citations. Note that our method is a training-free plug-and-play plugin that is capable of various LLMs. Experiments on various hallucination-related datasets show our framework outperforms state-of-the-art methods in both hallucination detection and response regeneration on three benchmarks. Our codes and dataset will be publicly available.
@@ 3.2 Retrieval Augmentation Module
Query: For the response $R$, it will be segmented into several claims if necessary, resulting in $R=\{R_{1},R_{2},...,R_{n}\}$. Here, we adhere to previous work Chen et al. (2023) and employ a heuristic algorithm for segmentation. Utilizing the NLTK Bird et al. (2009) sentence tokenizer, we split $R$ to obtain reasonable results that align with user reading habits to get the claims $R_{i}$. Then, $R_{i}$ is adopted as the query one by one.
@@
Where $e(·)$ is the SimCSE BERT encoder, $d_{j}$ is a candidate document in the corpus. As more documents need more calculation in further modules, the top-$k$ retrieved documents with higher similarity are selected to form the reference document set $D_{i}=\{d_{i,1},d_{i,2},...,d_{i,k}\}$. We add an extra threshold $t$ to filter out the retrieved documents that have low cosine similarity. Apart from the top-$1$ document, if the Sim($R_{i},d_{j}$) < $t$, $d_{j}$ will not be included in $D_{i}$. These documents are subsequently concatenated to construct the final retrieved content $D_{i}$ for further calculation.
@@ 3.3 Citation Generation Module
Although numerous models Honovich et al. (2022); Raffel et al. (2020) are capable of the NLI method, our CEG framework seeks to fully leverage the language comprehension capabilities of LLMs. Therefore, we prefer to utilize an LLM, equipped with predefined prompts, to serve as the NLI method. An illustrative prompt is provided below:
@@ 4.3 FELM Dataset
Baselines: Following settings in FELM, we adopt four strategies with ChatGPT, GPT-4, and Vicuna-33B as the backbone LLM Zheng et al. (2023): 1) Vanilla prompts. 2) Prompts augmented with Chain-of-Thought (CoT) reasoning Kojima et al. (2022). 3) Prompts augmented with hyperlinks to reference documents and 4) Prompts augmented with human-annotated reference documents Chen et al. (2023). Experiments are conducted at the individual response level. The metrics are accuracy of nonfactual, factual, and balanced.
4.4 HaluEval Dataset
@@
[…]
@@ 4.5 WikiRetr Datasets
WikiRetr datasets are designed to conduct further analyses on our CEG framework, which is created based on the October 20, 2023 snapshot of Wikipedia. We randomly select 1,000 passages, and apply text-davinci-003 and GPT-4 to rewrite them as new claims, separately. So that each rewritten claim is accompanied by an original passage. The constructed datasets are named WikiRetr-GPT3 and WikiRetr-GPT4, respectively.
@@ 5.1 Performance on Hallucination Detection
Experimental results in FELM dataset are summarized in Table 2. Firstly, our model achieves the best result in balanced accuracy with GPT-4, indicating its effectiveness. Most baseline models are biased in classifying a single type. Then, CEG with ChatGPT beats other ChatGPT baselines, showing the flexibility of our model. Thirdly, CEG outperforms all pre-retrieval baselines, which shows the strength of the proposed post-hoc segment-level retrieval module in hallucination detection.
@@ 5.3.1 Ablation Study
Table 4: Ablation results of CEG on the subset of FELM. ‘RA’ means the Retrieval Augmentation module.
We conduct ablation experiments on the FELM dataset, where the retrieval augmentation ($k$ = 4) and the document selection threshold are involved (threshold = 0.5). As shown in Table 4, the retrieval augmentation module plays an important role in providing better results on different backbone LLMs (ChatGPT and GPT-4). Furthermore, the threshold of retrieved documents is necessary, which can filter out irrelevant documents in citation generation. Thus, all designed modules contribute to improvements in the CEG framework.
@@ 6 Conclusion
Our study has several limitations: 1) Restricted retriever and corpus: In our experiments, we do not employ a fine-tuned specific retriever for post-hoc methods, and our method utilizes only the Wikipedia corpus, limiting the applicability of our framework to general knowledge-based question-answering scenarios but only demonstrate the effectiveness of our model. 2) Our experiments are conducted on existing benchmarks and manual annotations, where we do not propose new QA datasets for the verification of regeneration performance. 3) The adopted NLI method in the citation generation module inherently relies on the LLM’s world knowledge. More powerful NLI methods can contribute to better performance.
@@ Appendix A Prompts Used in Our Experiments
The prompt to regenerate a new response on the HaluEval dataset is presented in Table 13.
@@
The NLI prompt employed by GPT series models in experiments on WikiRetr datasets is presented above. The agreement rates with human annotaters of GPT models in Table 5 and the precision of GPT models in Table 6 are based on this prompt.
@@ Appendix C Case Studies on HaluEval Dataset
There are two cases in evaluation on HaluEval dataset, presented in Table 14 and Table 15, respectively.
@@
User:
@@
#Choice2#: Answer 1: The Lemon Drop Kid

### v4

@@ Abstract
Large language models (LLMs) exhibit powerful general intelligence across diverse scenarios, including their integration into chatbots. However, a vital challenge of LLM-based chatbots is that they may produce hallucinated content in responses, which significantly limits their applicability. Various efforts have been made to alleviate hallucination, such as retrieval augmented generation and reinforcement learning with human feedback, but most of them require additional training and data annotation. In this paper, we propose a novel post-hoc Citation-Enhanced Generation (CEG) approach combined with retrieval argumentation. Unlike previous studies that focus on preventing hallucinations during generation, our method addresses this issue in a post-hoc way. It incorporates a retrieval module to search for supporting documents relevant to the generated content, and employs a natural language inference-based citation generation module. Once the statements in the generated content lack of reference, our model can regenerate responses until all statements are supported by citations. Note that our method is a training-free plug-and-play plugin that is capable of various LLMs. Experiments on various hallucination-related datasets show our framework outperforms state-of-the-art methods in both hallucination detection and response regeneration on three benchmarks. Our code and datasets can be found at https://github.com/Tsinghua-dhy/CEG.
footnotetext: $\dagger$ Weizhi Ma (mawz@tsinghua.edu.cn) and Yang Liu (liuyang2011@tsinghua.edu.cn) are corresponding authors.
@@ 3.2 Retrieval Augmentation Module
Query: For the response $R$, it will be segmented into several claims if necessary, resulting in $R=\{R_{1},R_{2},...,R_{n}\}$. Here, we adhere to previous work Chen et al. (2023) and employ a heuristic algorithm for segmentation using the NLTK Bird et al. (2009) sentence tokenizer. The NLTK sentence tokenizer is a well-performing and widely used  Chen et al. (2023); Bird et al. (2009); Liu et al. (2023) sentence tokenizer and generally segments text correctly in most cases. We split $R$ to obtain reasonable results that align with user reading habits to get the claims $R_{i}$. Then, $R_{i}$ is adopted as the query one by one.
@@
Where $e(·)$ is the SimCSE BERT encoder, $d_{j}$ is a candidate document in the corpus. As more documents need more calculation in further modules, the top-$k$ retrieved documents with higher similarity are selected to form the reference document set $D_{i}$. We add an extra threshold $t$ to filter out the retrieved documents that have low cosine similarity. Apart from the top-$1$ document, if the Sim($R_{i},d_{j}$) < $t$, $d_{j}$ will not be included in $D_{i}$. These documents are subsequently concatenated to construct the final retrieved content $D_{i}$ for further calculation.
@@ 3.3 Citation Generation Module
Although numerous models Honovich et al. (2022); Raffel et al. (2020) are capable of the NLI method, our CEG framework seeks to fully leverage the language comprehension capabilities of LLMs. Therefore, we prefer to utilize an LLM with predefined prompts to serve as the NLI method. An illustrative prompt is provided below:
@@ 4.3 FELM Dataset
Baselines: Following settings in FELM, we adopt four strategies with ChatGPT, GPT-4, and Vicuna-33B as the backbone LLM Zheng et al. (2023): 1) Vanilla prompts. 2) Prompts augmented with Chain-of-Thought (CoT) reasoning Kojima et al. (2022). 3) Prompts augmented with hyperlinks to reference documents and 4) Prompts augmented with human-annotated reference documents Chen et al. (2023). Experiments are conducted at the individual response level. Following the previous work Chen et al. (2023), we chose accuracy of nonfactual, factual, and balanced as final metrics to to facilitate comparison with previous works.
@@
HalluDetector $\mathit{C}_{\textit{M}}=14,\mathit{C}_{\textit{FA}}=24$ 82.42 57.01 70.54
$\mathit{C}_{\textit{M}}=28,\mathit{C}_{\textit{FA}}=96$ 86.45 61.96 74.82
Focus $LLaMA-30B_{\textit{focus}}$ 89.79 65.69 73.64
$LLaMA-65B_{\textit{focus}}$ 89.94 64.90 74.08
@@ 4.5 WikiRetr Datasets
WikiRetr datasets are designed to conduct further analyses on our CEG framework, which is created based on the October 20, 2023 snapshot of Wikipedia. We randomly select 1,000 passages, and apply text-davinci-003 and GPT-4 to rewrite them as new claims, separately. So that each rewritten claim is accompanied by an original passage. The constructed datasets are named WikiRetr-GPT3 and WikiRetr-GPT4, respectively. Discussion about the reliability of WikiRetr datasets is provided in Appendix D.
@@ 5.1 Performance on Hallucination Detection
Experimental results in FELM dataset are summarized in Table 2. Firstly, our model achieves the best result in balanced accuracy with GPT-4, indicating its effectiveness. Most baseline models are biased in classifying a single type. Then, CEG with ChatGPT beats other ChatGPT baselines, showing the flexibility of our model. Thirdly, CEG outperforms all pre-retrieval baselines, which shows the strength of the proposed post-hoc segment-level retrieval module in hallucination detection. Finally, for Vicuna-33B, all methods exhibit some degree of decline compared to the Vallina method, indicating its limitations in general ability and using retrieved documents. However, our method shows the smallest decline, especially compared to the manually labeled Doc baseline, our method outperforms by 2.2 points, proving the effectiveness of our finer-grained document utilization.
@@ 5.3.1 Ablation Study
Table 4: Ablation results of CEG on the Worldknowledge subset of FELM. ‘RA’ means the Retrieval Augmentation module.
We conduct ablation experiments on the FELM Worldknowledge dataset, where the retrieval augmentation ($k$ = 4) and the document selection threshold are involved (threshold = 0.5). As shown in Table 4, the retrieval augmentation module plays an important role in providing better results on different backbone LLMs (ChatGPT and GPT-4). Furthermore, the threshold of retrieved documents is necessary, which can filter out irrelevant documents in citation generation. Thus, all designed modules contribute to improvements in the CEG framework.
@@ 6 Conclusion
Our study has several limitations: 1) Restricted retriever and corpus: In our experiments, we do not employ a fine-tuned specific retriever for post-hoc methods, and our method utilizes only the Wikipedia corpus, limiting the applicability of our framework to general knowledge-based question-answering scenarios but only demonstrate the effectiveness of our model. 2) Our experiments are conducted on existing benchmarks and manual annotations, where we do not propose new QA datasets for the verification of regeneration performance. 3) The adopted NLI method in the citation generation module inherently relies on the LLM’s world knowledge. More powerful NLI methods can contribute to better performance. 4) Prompting to regenerate and Using NLI technology to generate citations both incur API cost. API costs incurred for conducting our method and creating the WikiRetr datasets are shown in Table 11.
Acknowledgement
This work is supported by the National Natural Science Foundation of China (No. 62276152, 61925601, 62372260). We appreciate all the reviewers for their insightful suggestions.
@@ Appendix A Prompts Used in Our Experiments
The prompts for the baseline Vanilla, CoT, and Pre-Retrieval methods used in the HaluEval dataset are shown in Table 14, Table 15, and Table 16. The prompt to regenerate a new response on the HaluEval dataset is presented in Table 17.
@@
The NLI prompt employed by GPT series models in experiments on WikiRetr datasets is presented above. The agreement rates with human annotaters of GPT models in Table 5 and the precision of GPT models in Table 6 are based on this.
@@ Appendix C Case Studies on HaluEval Dataset
We conduct case studies on two cases in the evaluation of the HaluEval dataset, presented in Table 18 and Table 19, respectively.
Appendix D Discussion about the reliability of WikiRetr datasets
After constructing WikiRetr datasets, we randomly select 100 samples from each of the two datasets, and conduct manual annotation to assess whether the original passages can support the rewritten claims by three annotators. More specifically, each sample is initially annotated by two human annotators. In cases where there are discrepancies between the annotations provided by the two annotators, the final label is determined by consensus among three annotators.
Labeling results show that 90% and 94% of generated claims are supported by original documents. This consistency is exceptionally high. For example, in the FELM dataset, the average inter-annotator agreement for labels is 90.7%, and in the Pinocchio dataset Hu et al. (2023), the average label accuracy for sampled data is 92.7%, and the inter-annotator agreement is 85.6%.
Experiments GPT-3.5 GPT-4
Main result on FELM $\sim 700$ $\sim 700$
Main result on WikiBio GPT-3 $\sim 2000$
Main result on HaluEval $\sim 7700$
Variants of CEG on FELM $\sim 1400$ $\sim 1400$
Top-k ablation on FELM $\sim 6300$ $\sim 6300$
Top-k ablation on WikiBio GPT-3 $\sim 12000$
Creating WikiRetr datasets $\sim 1000$ $\sim 1000$
NLI experiments on WikiRetr datasets - $\sim 20000$
Annotation on WikiRetr datasets $\sim 400$ $\sim 400$
Table 11: API costs incurred for conducting our method and creating the WikiRetr datasets. We report the number of calls for different GPT models. For GPT-3.5, the total number of calls includes both GPT-3.5-Turbo-1106 and GPT-3.5-Turbo-Instruct. For GPT-4, the total number of calls includes both GPT-4-0613 and GPT-4-1106-preview.
@@ Appendix D Discussion about the reliability of WikiRetr datasets
Instruction: I want you act as an answer judge. Given a question, two answers, your objective is to select the best and correct answer without hallucination and nonfactual information. You should try your best to select the best and correct answer. If the two answers are the same, you can randomly choose one. If both answers are incorrect, choose the better one. You MUST select an answer from the provided two answers. Your response should be “Answer 1” or “Answer 2”.
#Question#: { Question }
#Answer 1#: { Right_answer }
#Answer 2#: { Hallucinated_answer }
Table 14: Prompt for the baseline Vanilla used in the HaluEval dataset.
[…]
#Question#: { Question }
#Answer 1#: { Right_answer }
#Answer 2#: { Hallucinated_answer }
Table 15: Prompt for the baseline CoT used in the HaluEval dataset.
[…]
#Question#: { Question }
#Answer 1#: { Right_answer }
#Answer 2#: { Hallucinated_answer }
Table 16: Prompt for the baseline Pre-Retrieval used in the HaluEval dataset.
User (Round 1):
@@
##: Answer 1: The Lemon Drop Kid